# Weak compactness cardinals for strong logics and subtlety properties of the class of ordinals

### v1

@@ 4. Separating the schemes
First, assume that (1) fails. Then Lemma 2.3 shows that there is a set-theoretic formula $\varphi(v_{0},v_{1})$ and a subset $z$ of $\kappa$ with the property that there exists a cardinal $\kappa<\vartheta\in C^{(n)}$such that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\vartheta)$ and there is no cardinal $\bar{\theta}\in C^{(n)}$ such that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\theta})$ for some cardinal $\bar{\kappa}<\bar{\theta}$. Since $\kappa$ is weakly $C^{(n)}$-shrewd, we can find a cardinal $\bar{\theta}\in C^{(n)}$ and a cardinal $\bar{\kappa}<\min(\kappa,\bar{\theta})$ such that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\theta})$. Now, define $\varepsilon$ to be the least element of $C^{(n)}$ above $\bar{\kappa}$ with the property that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\varepsilon)$. We then know that $\varepsilon\geq\kappa$ and it is easy to see that the set $\{\varepsilon\}$ is definable by a $\Sigma_{n+1}$-formula with parameters $\bar{\kappa},z\cap\bar{\kappa}\in{\rm{H}}(\kappa)$. This shows that (3) fails and hence we know that (3) implies (1). This completes the proof of the lemma, because (2) obviously implies (3) and Proposition 2.2.(3) shows that (1) implies (2). ∎
@@
The next lemma provides us with a tool to derive non-trivial consistency strength from the existence of $C^{(n)}$-weakly shrewd cardinals that are not $C^{(n)}$-strongly unfoldable. An intriguing aspect of this result is the fact that it establishes the existence of a subtle cardinal not only in some specific canonical inner model (e.g., in the constructible universe ${\rm{L}}$), but in all canonical inner models (e.g., all submodels of the class ${\rm{HOD}}$).
@@
Fix a $\Sigma_{n+1}$-formula $\varphi(v_{0},v_{1})$ and an element $y$ with the property that $\delta$ is the unique set $x$ such that $\varphi(x,y)$ holds. Pick a cardinal $\kappa<\theta\in C^{(n+1)}$ and use Lemma 3.3 find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\bar{\kappa}<\min(\kappa,\bar{\theta})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $y,\delta\in{{\rm{ran}}(j)}$. Then $y\in{\rm{H}}(\bar{\kappa})$ with $j(y)=y$. Pick $\varepsilon\in X$ with $j(\varepsilon)=\delta$. Since $\varphi(\delta,y)$ holds in ${\rm{H}}(\theta)$, we then know that $\varphi(\varepsilon,y)$ holds in ${\rm{H}}(\bar{\theta})$ and the fact $\bar{\theta}$ is an element of $C^{(n)}$ implies that $\varphi(\varepsilon,y)$ also holds in ${\rm{V}}$. This shows that $\delta=\varepsilon$ and hence $\delta>\kappa$. Next, note that the set $\{|{\delta}|\}$ is $\Sigma_{2}$-definable from the parameter $\delta$ and, since $|{\delta}|\geq\kappa$, the minimality of $\delta$ implies that $\delta$ is a cardinal. Finally, since the set $\{{{\rm{cof}}(\delta)}\}$ is $\Sigma_{2}$-definable from the parameter $\delta$, the minimality of $\delta$ also ensures that either $\delta$ is regular or ${{\rm{cof}}(\delta)}<\kappa$ holds.
@@
Assume, towards a contradiction, that $\kappa$ is not an element of $C$ and set $\alpha=\max(C\cap\kappa)<\kappa$ and $\beta=\min(C\setminus\kappa)=\min(C\setminus(\alpha+1))<\delta$. Then $\beta$ is an ordinal greater than $\kappa$ with the property that the set $\{\beta\}$ is definable by a $\Sigma_{n+1}$-formula with parameters $\alpha,z\in{\rm{H}}(\kappa)$ and hence it follows that $\beta\geq\delta$, a contradiction. ∎
@@
is consistent and there is a sentence $\phi$ in the language of set theory with the property that the statements (1) and (2) listed in the corollary hold. By combining Theorem 1.7 with Theorem 2.4, it follows that the theory
@@
is also consistent. Define $\psi$ to be the conjunction of $\phi$ with the statement that only boundedly many cardinals are subtle in ${\rm{HOD}}$. Then (1) together with the above observations ensures that ${\rm{ZFC}}+\psi$ is consistent. Moreover, by (2), we can apply Theorem 1.7 to conclude that
@@
holds, contradicting Proposition 1.9. ∎
@@
In order to expand our analysis of the given class principles, we now explore a phenomenon already unveiled by [13, Theorem 1.11] showing that the existence of weakly shrewd that are not strongly unfoldable yields reflection properties for statements of arbitrary high complexities:
@@
Pick $y\in{\rm{H}}(\kappa)$ with the property that the set $\{\delta\}$ is definable by a $\Sigma_{m+1}$-formula with parameter $y$. Since $\kappa$ is a limit cardinal, we can find a cardinal $\alpha<\rho<\kappa$ with $y\in{\rm{H}}(\rho)$. Moreover, since $n>m$, it follows that the interval $(\rho,\delta)$ does not contain an element of $C^{(n+1)}$. Now, assume, towards a contradiction, the interval $(\rho,\delta)$ does not contain a $C^{(n)}$-weakly shrewd cardinal that is not $C^{(n)}$-strongly unfoldable. The above observation then allow us to use Lemma 4.1 to conclude that the interval $(\rho,\delta)$ does not contain $C^{(n)}$-weakly shrewd cardinals. This shows that for every cardinal $\mu$ in the interval $(\rho,\delta)$, there exists a set-theoretic formula $\varphi(v_{0},v_{1})$, a cardinal $\mu<\theta\in C^{(n)}$ and a subset $z$ of $\mu$ such that $\varphi(z,\mu)$ holds in ${\rm{H}}(\theta)$ and $\varphi(z\cap\bar{\mu},\bar{\mu})$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\bar{\mu}<\min(\mu,\bar{\theta})$. We can then find a cardinal $\delta<\zeta\in C^{(1)}$ with the property that, in ${\rm{H}}(\zeta)$, for every cardinal $\mu$ in the interval $(\rho,\delta)$, there exists a set-theoretic formula $\varphi(v_{0},v_{1})$, a cardinal $\mu<\theta\in C^{(n)}$ and a subset $z$ of $\mu$ such that $\varphi(z,\mu)$ holds in ${\rm{H}}(\theta)$ and $\varphi(z\cap\bar{\mu},\bar{\mu})$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\bar{\mu}<\min(\mu,\bar{\theta})$. Let $\eta$ be the least cardinal with this property. Then the set $\{\eta\}$ is definable by a $\Sigma_{2}$-formula with parameters $\delta$ and $y$. We then directly know that the set $\{\eta\}$ is also definable by a $\Sigma_{m+1}$-formula with parameter $y$.
@@
Now, pick $\kappa<\vartheta\in C^{(m+1)}$ and use Lemma 3.3 to find $\bar{\vartheta}\in C^{(m)}$, a cardinal $\bar{\kappa}<\min(\kappa,\bar{\vartheta})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\vartheta})$ with $\bar{\kappa}+1\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\vartheta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $y,\rho\in{{\rm{ran}}(j)}$. Then $y\in X$ with $j(y)=y$. Moreover, we know that $\rho<\bar{\kappa}$ and therefore $j(\rho)=\rho$. In combination with the facts that all $\Sigma_{m+1}$-statements are upwards absolute from $X$ to ${\rm{V}}$, this shows that $\delta,\eta\in X$ with $j(\delta)=\delta$ and $j(\eta)=\eta$. Since $\bar{\kappa}$ is a cardinal in the interval $(\rho,\delta)$ in ${\rm{H}}(\eta)$, we can now find a set-theoretic formula $\varphi(v_{0},v_{1})$, an ordinal $\bar{\kappa}<\theta\in X\cap\eta$ and $z\in{\mathcal{P}}({\bar{\kappa}})\cap X$ with the property that, in ${\rm{H}}(\eta)$, the ordinal $\theta$ is a cardinal in $C^{(n)}$ such that $\varphi(z,\bar{\kappa})$ holds in ${\rm{H}}(\theta)$ and $\varphi(z\cap\mu,\mu)$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\mu<\min(\bar{\kappa},\bar{\theta})$. The fact that ${\rm{H}}(\eta)\in X$ with $j({\rm{H}}(\eta))={\rm{H}}(\eta)$ then allows us to use the elementarity of $j$ to conclude that, in ${\rm{H}}(\eta)$, the ordinal $j(\theta)$ is a cardinal in $C^{(n)}$ with the property that $\varphi(j(z),\kappa)$ holds in ${\rm{H}}(j(\theta))$ and $\varphi(j(z)\cap\mu,\mu)$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\mu<\min(\kappa,\bar{\theta})$. But, this yields a contradiction, because $\theta$ is an element of $C^{(n)}$ in ${\rm{H}}(\eta)$, $\bar{\kappa}<\min(\kappa,\theta)$ and $j(z)\cap\bar{\kappa}=z$. ∎
@@
Fix an ordinal $\alpha>\omega$ and pick $\delta\in S$ with $R(\delta)>\alpha$. Set $S_{0}=S\cap\delta$ and let ${\mathbb{Q}}$ denote the Easton support product of all partial orders of the form ${\rm{Add}}({R(\gamma)},{\gamma})$ with $\gamma\in S_{0}$, as constructed in ${\rm{V}}$. In addition, let ${\mathbb{R}}$ denote the partial order ${\rm{Add}}({R(\delta)},{\delta})$ as constructed in ${\rm{V}}$. A standard factor analysis of the class partial order ${\mathbb{P}}$ (again, as in [10, pp. 235–236]) now allows us to find an inner model $M$ of ${\rm{V}}[G]$ such that ${\rm{V}}\subseteq M$, $({}^{\delta}{\rm{V}})^{M}\subseteq{\rm{V}}$ and there is are $G_{0},G_{1}\in{\rm{V}}[G]$ such that $G_{0}\times G_{1}$ is $({\mathbb{Q}}\times{\mathbb{R}})$-generic over $M$ with ${\rm{V}}[G]=M[G_{0},G_{1}]$. Then $\delta$ is a subtle cardinal in $M$ and the definition of $R$ in ${\rm{V}}$ ensures that $\delta$ is the minimal subtle cardinal above $R(\delta)$ in $M$. In particular, it follows that the set $\{\delta\}$ is definable by a $\Sigma_{2}$-formula with parameter $R(\delta)$ in $M$.

### v3

@@ 4. Separating the schemes
First, assume that (1) fails and (3) holds. Given $\alpha<\kappa$, the fact that the set $\{|{{\rm{V}}_{\alpha}}|\}$ is definable by a $\Sigma_{2}$-formula with parameter $\alpha$ allows us to use our assumption to conclude that $|{{\rm{V}}_{\alpha}}|<\kappa$ holds. In particular, we know that the set ${\rm{V}}_{\kappa}$ has cardinality $\kappa$ in this setting. Since (1) fails, Lemma 2.3 shows that there is a set-theoretic formula $\varphi(v_{0},v_{1})$, a cardinal $\kappa<\theta\in C^{(n)}$ and a subset $A$ of ${\rm{V}}_{\kappa}$ such that $\varphi(A,\kappa)$ holds in ${\rm{H}}(\theta)$ and there is no cardinal $\bar{\theta}\in C^{(n)}\cap\kappa$ such that $\varphi(A\cap{\rm{V}}_{\bar{\kappa}},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\theta})$ for some cardinal $\bar{\kappa}<\bar{\theta}$. Using the fact that $\kappa$ is weakly inaccessible and the set ${\rm{V}}_{\kappa}$ has cardinality $\kappa$, we can now find a set-theoretic formula $\psi(v_{0},v_{1})$ and a subset $z$ of $\kappa$ with the property that $\psi(z,\kappa)$ holds in ${\rm{H}}(\theta)$ and there is no cardinal $\bar{\theta}\in C^{(n)}\cap\kappa$ such that $\psi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\theta})$ for some cardinal $\bar{\kappa}<\bar{\theta}$. Since $\kappa$ is $C^{(n)}$-weakly shrewd, we can find a cardinal $\bar{\theta}\in C^{(n)}$ and a cardinal $\bar{\kappa}<\min(\kappa,\bar{\theta})$ such that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\theta})$. Now, define $\varepsilon$ to be the least element of $C^{(n)}$ above $\bar{\kappa}$ with the property that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\varepsilon)$. We then know that $\varepsilon\geq\kappa$ and it is easy to see that the set $\{\varepsilon\}$ is definable by a $\Sigma_{n+1}$-formula with parameters $\bar{\kappa},z\cap\bar{\kappa}\in{\rm{H}}(\kappa)$. This contradicts (3) and we can conclude that (3) implies (1).
These computations complete the proof of the lemma, because (2) obviously implies (3) and Proposition 2.2.(3) shows that (1) implies (2). ∎
@@
The next lemma provides us with a tool to derive non-trivial consistency strength from the existence of $C^{(n)}$-weakly shrewd cardinals that are not $C^{(n)}$-strongly unfoldable. An interesting aspect of this result is the fact that, in contrast to most results that provide lower bounds for the consistency strength of certain assumptions, it establishes the existence of large cardinals not only in certain fine-structural inner models (e.g., in the constructible universe ${\rm{L}}$) but also in the inner model ${\rm{HOD}}$. Since the results of this paper show that the given assumption is compatible with the non-existence of strongly inaccessible cardinals, the following lemma reveals that the existence of weakly shrewd cardinals that are not strongly unfoldable has interesting and non-trivial consequences for the behavior of ordinal definable sets.
@@
Fix a $\Sigma_{n+1}$-formula $\varphi(v_{0},v_{1})$ and an element $y$ of ${\rm{H}}(\kappa)$ with the property that $\delta$ is the unique set $x$ such that $\varphi(x,y)$ holds. Pick a cardinal $\kappa<\theta\in C^{(n+1)}$ and use Lemma 3.3 to find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\bar{\kappa}<\min(\kappa,\bar{\theta})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $y,\delta\in{{\rm{ran}}(j)}$. Then $y\in{\rm{H}}(\bar{\kappa})$ with $j(y)=y$. Pick $\varepsilon\in X$ with $j(\varepsilon)=\delta$. Since $\varphi(\delta,y)$ holds in ${\rm{H}}(\theta)$, we then know that $\varphi(\varepsilon,y)$ holds in ${\rm{H}}(\bar{\theta})$ and the fact that $\bar{\theta}$ is an element of $C^{(n)}$ implies that $\varphi(\varepsilon,y)$ also holds in ${\rm{V}}$. This shows that $\delta=\varepsilon$ and hence $\delta>\kappa$. Next, note that the set $\{|{\delta}|\}$ is $\Sigma_{2}$-definable from the parameter $\delta$ and, since $|{\delta}|\geq\kappa$, the minimality of $\delta$ implies that $\delta$ is a cardinal. Finally, since the set $\{{{\rm{cof}}(\delta)}\}$ is $\Sigma_{2}$-definable from the parameter $\delta$, the minimality of $\delta$ also ensures that either $\delta$ is regular or ${{\rm{cof}}(\delta)}<\kappa$ holds.
@@
Assume, towards a contradiction, that $\kappa$ is not an element of $C$ and set $\alpha=\max(C\cap\kappa)<\kappa$ and $\beta=\min(C\setminus\kappa)=\min(C\setminus(\alpha+1))<\delta$. Then $\beta$ is an ordinal greater than $\kappa$ with the property that the set $\{\beta\}$ is definable by a $\Sigma_{n+1}$-formula with parameters $\alpha,y\in{\rm{H}}(\kappa)$ and hence it follows that $\beta\geq\delta$, a contradiction. ∎
@@
is consistent and there is a sentence $\phi$ in the language of set theory with the property that the statements (1) and (2) listed in the corollary hold. Define $\psi$ to be the conjunction of $\phi$ with the statement that no ordinal is a subtle cardinal in ${\rm{HOD}}$.
Claim.
The theory ${\rm{ZFC}}+\psi$ is consistent.
Proof of the Claim.
By our assumption and Corollary 1.8, we know that the theory
ZFC+Ord is essentially ${\rm{ZFC}}+{\text{``}\hskip 1.29167pt{\textit{${\rm{Ord}}$ is essentially subtle}}\hskip 1.29167pt\text{''}}$ ”
is consistent and, since subtle cardinals yield set-sized models of this theory, it follows that the theory
@@
is also consistent. By (1), this theory proves $\psi$. ∎
By (2), we can now apply Theorem 1.7 to conclude that
@@
holds. In this situation, Proposition 1.9 shows that ${\rm{ZFC}}+\psi$ is inconsistent, contradicting the above claim. ∎
@@
In order to expand our analysis of the given class principles, we now explore a phenomenon already unveiled by [13, Theorem 1.11] showing that the existence of weakly shrewd cardinals that are not strongly unfoldable yields reflection properties for statements of arbitrary high complexities. The following lemma strengthens the result from [13] by both weakening the assumption used and establishing the existence of large cardinals that induce the given reflection properties:
@@
Pick $y\in{\rm{H}}(\kappa)$ with the property that the set $\{\delta\}$ is definable by a $\Sigma_{m+1}$-formula with parameter $y$. Since $\kappa$ is a limit cardinal, we can find a cardinal $\alpha<\rho<\kappa$ with $y\in{\rm{H}}(\rho)$. Moreover, since $n>m$, it follows that the interval $(\rho,\delta)$ does not contain an element of $C^{(n+1)}$. Now, assume, towards a contradiction, the interval $(\rho,\delta)$ does not contain a $C^{(n)}$-weakly shrewd cardinal that is not $C^{(n)}$-strongly unfoldable. The above observation then allow us to use Lemma 4.1 to conclude that the interval $(\rho,\delta)$ does not contain $C^{(n)}$-weakly shrewd cardinals. This shows that for every cardinal $\mu$ in the interval $(\rho,\delta)$, there exists a set-theoretic formula $\varphi(v_{0},v_{1})$, a cardinal $\mu<\theta\in C^{(n)}$ and a subset $z$ of $\mu$ such that $\varphi(z,\mu)$ holds in ${\rm{H}}(\theta)$ and $\varphi(z\cap\bar{\mu},\bar{\mu})$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\bar{\mu}<\min(\mu,\bar{\theta})$. We can then find a cardinal $\delta<\zeta\in C^{(1)}$ with the property that, in ${\rm{H}}(\zeta)$, for every cardinal $\mu$ in the interval $(\rho,\delta)$, there exists a set-theoretic formula $\varphi(v_{0},v_{1})$, a cardinal $\mu<\theta\in C^{(n)}$ and a subset $z$ of $\mu$ such that $\varphi(z,\mu)$ holds in ${\rm{H}}(\theta)$ and $\varphi(z\cap\bar{\mu},\bar{\mu})$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\bar{\mu}<\min(\mu,\bar{\theta})$. Let $\eta$ be the least cardinal with this property. Then the set $\{\eta\}$ is definable by a $\Sigma_{2}$-formula with parameters $\delta$, $\rho$ and $y$. We then directly know that the set $\{\eta\}$ is also definable by a $\Sigma_{m+1}$-formula with parameters $\rho$ and $y$.
@@
Now, pick $\kappa<\vartheta\in C^{(m+1)}$ and use Lemma 3.3 to find $\bar{\vartheta}\in C^{(m)}$, a cardinal $\bar{\kappa}<\min(\kappa,\bar{\vartheta})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\vartheta})$ with $\bar{\kappa}+1\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\vartheta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $\rho,y\in{{\rm{ran}}(j)}$. Then $y\in X$ with $j(y)=y$. Moreover, we know that $\rho<\bar{\kappa}$ and therefore $j(\rho)=\rho$. In combination with the fact that all $\Sigma_{m+1}$-statements are upwards absolute from $X$ to ${\rm{V}}$, this shows that $\delta,\eta\in X$ with $j(\delta)=\delta$ and $j(\eta)=\eta$. Since $\bar{\kappa}$ is a cardinal in the interval $(\rho,\delta)$ in ${\rm{H}}(\eta)$, we can now find a set-theoretic formula $\varphi(v_{0},v_{1})$, an ordinal $\bar{\kappa}<\theta\in X\cap\eta$ and $z\in{\mathcal{P}}({\bar{\kappa}})\cap X$ with the property that, in ${\rm{H}}(\eta)$, the ordinal $\theta$ is a cardinal in $C^{(n)}$ such that $\varphi(z,\bar{\kappa})$ holds in ${\rm{H}}(\theta)$ and $\varphi(z\cap\mu,\mu)$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\mu<\min(\bar{\kappa},\bar{\theta})$. The fact that ${\rm{H}}(\eta)\in X$ with $j({\rm{H}}(\eta))={\rm{H}}(\eta)$ then allows us to use the elementarity of $j$ to conclude that, in ${\rm{H}}(\eta)$, the ordinal $j(\theta)$ is a cardinal in $C^{(n)}$ with the property that $\varphi(j(z),\kappa)$ holds in ${\rm{H}}(j(\theta))$ and $\varphi(j(z)\cap\mu,\mu)$ does not hold in ${\rm{H}}(\bar{\theta})$ for all $\bar{\theta}\in C^{(n)}$ and all cardinals $\mu<\min(\kappa,\bar{\theta})$. But, this yields a contradiction, because $\theta$ is an element of $C^{(n)}$ in ${\rm{H}}(\eta)$, $\bar{\kappa}<\min(\kappa,\theta)$ and $j(z)\cap\bar{\kappa}=z$. ∎
@@
Fix an ordinal $\alpha>\omega$ and pick $\delta\in S$ with $R(\delta)>\alpha$. Set $S_{0}=S\cap\delta$ and let ${\mathbb{Q}}$ denote the Easton support product of all partial orders of the form ${\rm{Add}}({R(\gamma)},{\gamma})$ with $\gamma\in S_{0}$, as constructed in ${\rm{V}}$. In addition, let ${\mathbb{R}}$ denote the partial order ${\rm{Add}}({R(\delta)},{\delta})$ as constructed in ${\rm{V}}$. A standard factor analysis of the class partial order ${\mathbb{P}}$ (again, as in [10, pp. 235–236]) now allows us to find an inner model $M$ of ${\rm{V}}[G]$ such that ${\rm{V}}\subseteq M$, $({}^{\delta}{\rm{V}})^{M}\subseteq{\rm{V}}$ and there are $G_{0},G_{1}\in{\rm{V}}[G]$ such that $G_{0}\times G_{1}$ is $({\mathbb{Q}}\times{\mathbb{R}})$-generic over $M$ with ${\rm{V}}[G]=M[G_{0},G_{1}]$. Then $\delta$ is a subtle cardinal in $M$ and the definition of $R$ in ${\rm{V}}$ ensures that $\delta$ is the minimal subtle cardinal above $R(\delta)$ in $M$. In particular, it follows that the set $\{\delta\}$ is definable by a $\Sigma_{2}$-formula with parameter $R(\delta)$ in $M$.